OpenStudy (anonymous):

Find the derivative ln(ln7x) using chain rule

OpenStudy (anonymous):

First let \(u = \ln(7x)\) So we want to find \[ \frac{d}{dx} \ln(\ln(7x))= \frac{d}{dx}\ln(u) = \frac{d}{du}[\ln(u)]\frac{du}{dx} = \frac{1}{u}\frac{du}{dx} \]

OpenStudy (jdoe0001):

I'd say use the chain-rule, obviously =)

OpenStudy (anonymous):

Thank you, Wio

OpenStudy (anonymous):

@jwallen Well, that was enough to get you started. It wasn't the complete answer. I hope you figured the rest out on your own though.

OpenStudy (anonymous):

I think i plug in u in du/dx=1/u du/dx